# A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance

### v4

@@ 6 Conclusion
In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance. An inference attack, whereby an adversary estimates the private state, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level. We develop computational methods based on interval analysis, and establish the theoretical properties of the measure. Furthermore, we propose an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method is demonstrated through a production-inventory case study.
@@ Appendix C Proof of Theorem 9
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
@@
$=\left\|A_{3}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
@@
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
@@ Appendix D Proof of Theorem 12
Since $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (28) and (31) with the constraints of difference sets (C). Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative. As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (63), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.

### v5

@@ 6 Conclusion
In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance. An inference attack, whereby an adversary estimates the private information, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level. We develop computational methods based on interval analysis, and establish the theoretical properties of the measure. Furthermore, we propose an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method is demonstrated through a production-inventory case study.
@@ Appendix C Proof of Theorem 9
$\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
@@
$=\left\|A_{3}\right\|_{1}\overline{\mathrm{Vol}}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
@@
$\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
@@ Appendix D Proof of Theorem 12
Since $\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (32) and (35) with the constraints of difference sets (C). Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative. As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (68), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\overline{\mathrm{Vol}}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.